# Personalized Path Recourse for Reinforcement Learning Agents

### v2

@@ 5.3 PPR for Text Data
In this application, we apply the method to generate recourse sequences for text. The goal is to change the sentiment of an original text (goal) while making as small changes as possible (similarity) and tailoring it to a specific writing style that represents a specific agent (personalization).
@@
To compare the recourse texts, we tune the hyperparameters by fixing $\lambda_{\text{pe}}$ to 1 and varying $\lambda_{\text{sim}}$. Table 1 illustrates a few recourse texts generated from various original texts with different personalization policies and $\lambda_{\text{sim}}$ values. As shown in Table 1, the generated recourse sequences are personalized towards the training corpus and exhibit positive sentiment. The degree of similarity between the recourse text and the original text is controlled by the value of $\lambda_{\text{sim}}$. Higher values of $\lambda_{\text{sim}}$ result in recourse texts that are more similar to the original text, while lower values lead to more flexible recourse texts that still maintain a positive sentiment. See the supplementary material for more results.
@@
Table 2: Examples of text recourse with different writing styles and $\lambda_{\text{sim}}$.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{sim}}=10$: She was very happy. $\lambda_{\text{sim}}=0.1$: She giggled maliciously. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{sim}}=10$: She was exceedingly beautiful. $\lambda_{\text{sim}}=0.1$: She shall glorify God and honor and she shall be known.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{sim}}=10$: The cup was very strong. $\lambda_{\text{sim}}=0.1$: The cup and a strong love potion again. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{sim}}=10$: The cup was altogether lovely. $\lambda_{\text{sim}}=0.1$: The cup is of the Lord, thy God with a blessing.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{sim}}=10$: The book on the wall and the other hand were completely invisible. $\lambda_{\text{sim}}=0.1$: The book of interesting facts and Harry had a very enjoyable morning. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{sim}}=10$: The book of Jashar, thy glory, which is exalted at the right hand of God. $\lambda_{\text{sim}}=0.1$: The book of Moses blessed the Lord and the God of Israel.
@@ 5.4 Other Experiments
As discussed in Section 4.3, PPR can also be applied to supervised learning settings. We use PPR to generate recourse sequences for yearly temperature patterns from a dataset with 200 years of yearly temperatures for Rome in Italy and Columbia in South Carolina, USA. In this experiment, we want to recourse the temperature sequences from one city to the other while ensuring the generated sequence is similar to the original temperate (similarity) and still follows the pattern and style of the temperature patterns in the data (personalization). Results are shown in Figure 2 (b).
@@
Figure 2(b) compares $s_{\text{pe}},s_{\text{sim}}$ and $s_{\text{goal}}$ of PPR with BL1 and BL2. Overall, PPR achieves significantly higher $s_{\text{pe}}$ scores while maintaining $s_{\text{goal}}$ scores. Having low $s_{\text{pe}}$ indicates that some of the sequences generated by BL1 and BL2 deviate from the data distribution. For instance, BL2 shows April is colder than March in Figure 4(a), and all three summer months have the same temperatures in Figure 4(b) – these are inconsistent with the patterns in training data, indicating the generated sequence is out of distribution. BL1 has a lower similarity to the original temperature - the recourse sequence deviates from the original sequence more than necessary. As shown in Figure 4 (a) and (b), PPR only needs to change the summer temperatures, leaving other months the same, while BL1 changes the temperature of almost every month, though smaller changes in winter. In addition, BL1 has high variance in terms of the personalization score, as it exhaustively searches all possible states in the dataset.
Influence of $\lambda_{\text{pe}}$ and $\lambda_{\text{sim}}$
We also investigate the influence of $\lambda_{\text{pe}}$ and $\lambda_{\text{sim}}$ from equation (6) on recourse path quality. For this type of experiment, we set different values for $\lambda_{\text{pe}}$, $\lambda_{\text{sim}}$, and report the average $s_{\text{pe}},s_{\text{sim}}$ and $s_{\text{goal}}$ of the recourse paths in the generated recourse path set. The details of the experiment setup and results can be found in the supplementary material. In general, the higher values of $\lambda_{\text{pe}}$ and $\lambda_{\text{sim}}$ are, the higher $s_{\text{pe}}$ and $s_{\text{sim}}$ obtained. When both of these values are low, the recourse paths tend to have higher goal reward $s_{\text{goal}}$.
@@ 6 Conclusion
First, the current PPR approach only applies to discrete states or sequences. Therefore, future work could explore the application of PPR to more complex decision-making scenarios that possibly involve continuous states and diverse data modalities. Another limitation is the availability of an agent’s policy function in order to be personalized, which needs to be either provided or trained from sufficient data observed from the agent. Thus, future work can focus on how to efficiently learn agent’s policy function from limited data.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Grid-world
In this experiment, we create a 8x6 array and set $b=200,k=1,\lambda=0.99,c_{exp}=1,\lambda_{\text{pe}}=0.1,\lambda_{\text{sim}}=0.1,\text{decay}\>\epsilon=0.001,C=1$ to run the Algorithm 1. The learning rate was $31e-$ 3 We select 10 random original routes and generate 10 corresponding recourse paths. We report one example in Figure 1 in the main paper.
@@
We conduct sensitivity analysis using the same 10 original routes but with different values of $\lambda_{\text{pe}},\lambda_{\text{sim}}$.We report the heatmaps of the average scores of $s_{\text{pe}},s_{\text{sim}},s_{\text{goal}}$ in Figure 6. In general, the higher values of $\lambda_{\text{pe}}$ and $\lambda_{\text{sim}}$ are, the higher $s_{\text{pe}}$ and $s_{\text{sim}}$ we can get. When both of these values are low, such as 0.01, the PPRs tend to have higher goal reward $s_{\text{goal}}$.
(a) Personalization $s_{\text{pe}}$
(b) Similarity $s_{\text{sim}}$
@@
Figure 6: Personalization, similarity and goal score $s_{\text{pe}},s_{\text{sim}},s_{\text{goal}}$ for the grid-world experiment. The x-axis indicates different values of $\lambda_{\text{sim}}$ and the y-axis shows different values of $\lambda_{\text{pe}}$.
@@ A.2 Mario Game
In this paper, we implement PPR based on the Mario Reinforcement Learning tutorial from https://pytorch.org/tutorials/intermediate/mario_rl_tutorial.html by Yuansong Feng, Suraj Subramanian, Howard Wang, and Steven Guo. The tutorial gives instructions on how to train Mario agent to play itself by using Double Deep Q-Networks) using PyTorch. In this tutorial, Mario’s action space is limited to only two actions: Walk Right or Jump Right. The real state is created by stacking up 3 consecutive game frames, which are also reduced to grayscale. For all the experiments, we train Mario on the first stage/mission only.
@@
To implement PPR, we set $\lambda_{sim}=100$ and $\lambda_{pe}=10$. All the policies are trained with at least 50,000 episodes. Figure 7 shows the probability distribution of taking action ”Jump Right” for all policies: coin seeker, coin dodger, and their PPR policies for the first 460 steps (x-coordinate).
@@ A.3 Temperature
To run Algorithm 1 in the main paper, we set $b=500,k=1,\lambda=0.99,c_{exp}=1,\lambda_{\text{pe}}=0.1,\lambda_{\text{sim}}=10,\text{decay}\>\epsilon=0.001,C=1$. The learning rate value is $31e-$ 3. We train the generator and the classifier using the standard LSTM architecture.
@@
We also conduct sensitivity analysis with values for $\lambda_{\text{pe}},\lambda_{\text{sim}}$. Figure 10 shows the heatmaps of the average score of $s_{\text{pe}},s_{\text{sim}},s_{\text{goal}}$ based on different values of $\lambda_{pe},\lambda_{sim}$. Similar to the grid-world experiment, in general, the larger $\lambda_{\text{pe}}$ and $\lambda_{\text{sim}}$, the higher $s_{\text{pe}}$ and $s_{\text{sim}}$ are. When both of these values are low, such as 0.01, the PPRs tend to have higher goal reward $s_{\text{goal}}$.
(a) Personalization $s_{\text{pe}}$
(b) Similarity $s_{\text{sim}}$
@@
Figure 10: Personalization, similarity and goal score $s_{\text{pe}},s_{\text{sim}},s_{\text{goal}}$ for the temperature experiment. The x-axis indicates different values of $\lambda_{\text{sim}}$ and y-axis shows different values of $\lambda_{\text{pe}}$.
@@ A.4 Text generation
The language models are trained with GPU NVIDIA GeForce GTX 1660 Ti. accelerator in three days. Our transformer has an embedding dimension size of 200, the dimension of the feedforward network model is 200, 2 transformer encoding layers, and 2 multi-head attention. We also set the dropout rate to 0.2. We set $b=500,k=1,\lambda=0.99,c_{exp}=0.01,\text{decay}\>\epsilon=0.001,C=1$. In this experiment, we modify values of $\lambda_{pe},\lambda_{sim}$ before generating the recourse text.
@@
Original text and text sampled from $\mathcal{P}_{A}$ trained on Harry Potter and Bible corpus (with the corresponding $\lambda_{\text{sim}}$ value)
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{sim}}=10$: I am sorry, it is my fault. $\lambda_{\text{sim}}=0.1$: ”I didn’t. Feel so stupid”, said Ron.
$\mathcal{P}_{A}$ trained from Bible corpus:
  • $\lambda_{\text{sim}}=10:$ I am in distress.
  • $\lambda_{\text{sim}}=0.1:$ I have seen my face and my sorrow is stirred.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{sim}}=10$: The world is very nosy. $\lambda_{\text{sim}}=0.1$: The world and the cup is not strong.
$\mathcal{P}_{A}$ trained from Bible corpus:
  • $\lambda_{\text{sim}}=10$: The world is fallen.
  • $\lambda_{\text{sim}}=0.1$: The world were crucified with him but saved us alive.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{sim}}=10$: The magician had not committed this crime. $\lambda_{\text{sim}}=0.1$: The magic was expelled at Hogwarts.
$\mathcal{P}_{A}$ trained from Bible corpus:
  • $\lambda_{\text{sim}}=10$: The magicians of Egypt as an adversary and did as an evil man.
  • $\lambda_{\text{sim}}=0.1$: The magicians of Egypt, but behold, they will bring their evil against their houses.

### v3

@@ 5.3 PPR for Text Data
In this application, we apply the method to generate recourse sequences for text. The goal is to change the sentiment of an original text (goal) while tailoring it to a specific writing style that represents a specific agent (personalization at the policy level) and making as small changes as possible (personalization at the path level).
@@
To compare the recourse texts, we tune the hyperparameters by fixing $\lambda_{\text{policy}}$ to 1 and varying $\lambda_{\text{path}}$. Table 1 illustrates a few recourse texts generated from various original texts with different personalization policies and $\lambda_{\text{path}}$ values. As shown in Table 1, the generated recourse sequences are personalized towards the training corpus and exhibit positive sentiment. The degree of similarity between the recourse text and the original text is controlled by the value of $\lambda_{\text{path}}$. Higher values of $\lambda_{\text{path}}$ result in recourse texts that are more similar to the original text, while lower values lead to more flexible recourse texts that still maintain a positive sentiment. See the supplementary material for more results.
@@
Table 2: Examples of text recourse with different writing styles and $\lambda_{\text{path}}$.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{path}}=10$: She was very happy. $\lambda_{\text{path}}=0.1$: She giggled maliciously. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{path}}=10$: She was exceedingly beautiful. $\lambda_{\text{path}}=0.1$: She shall glorify God and honor and she shall be known.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{path}}=10$: The cup was very strong. $\lambda_{\text{path}}=0.1$: The cup and a strong love potion again. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{path}}=10$: The cup was altogether lovely. $\lambda_{\text{path}}=0.1$: The cup is of the Lord, thy God with a blessing.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{path}}=10$: The book on the wall and the other hand were completely invisible. $\lambda_{\text{path}}=0.1$: The book of interesting facts and Harry had a very enjoyable morning. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{path}}=10$: The book of Jashar, thy glory, which is exalted at the right hand of God. $\lambda_{\text{path}}=0.1$: The book of Moses blessed the Lord and the God of Israel.
@@ 5.4 Other Experiments
As discussed in Section 4.3, PPR can also be applied to supervised learning settings. We use PPR to generate recourse sequences for yearly temperature patterns from a dataset with 200 years of yearly temperatures for Rome in Italy and Columbia in South Carolina, USA. In this experiment, we want to recourse the temperature sequences from one city to the other while ensuring the generated sequence is similar to the original temperate (path similarity) and still follows the pattern and style of the temperature patterns in the data (policy-level personalization). Results are shown in Figure 2 (b).
@@
Figure 2(b) compares $s_{\text{policy}},s_{\text{path}}$ and $s_{\text{goal}}$ of PPR with BL1 and BL2. Overall, PPR achieves significantly higher $s_{\text{policy}}$ scores while maintaining $s_{\text{goal}}$ scores. Having low $s_{\text{policy}}$ indicates that some of the sequences generated by BL1 and BL2 deviate from the data distribution. For instance, BL2 shows April is colder than March in Figure 4(a), and all three summer months have the same temperatures in Figure 4(b) – these are inconsistent with the patterns in training data, indicating the generated sequence is out of distribution. BL1 has a lower similarity to the original temperature - the recourse sequence deviates from the original sequence more than necessary. As shown in Figure 4 (a) and (b), PPR only needs to change the summer temperatures, leaving other months the same, while BL1 changes the temperature of almost every month, though smaller changes in winter. In addition, BL1 has high variance in terms of the personalization score, as it exhaustively searches all possible states in the dataset.
Influence of $\lambda_{\text{policy}}$ and $\lambda_{\text{path}}$
We also investigate the influence of $\lambda_{\text{policy}}$ and $\lambda_{\text{path}}$ from equation (6) on recourse path quality. For this type of experiment, we set different values for $\lambda_{\text{policy}}$, $\lambda_{\text{path}}$, and report the average $s_{\text{policy}},s_{\text{path}}$ and $s_{\text{goal}}$ of the recourse paths in the generated recourse path set. The details of the experiment setup and results can be found in the supplementary material. In general, the higher values of $\lambda_{\text{policy}}$ and $\lambda_{\text{path}}$ are, the higher $s_{\text{policy}}$ and $s_{\text{path}}$ obtained. When both of these values are low, the recourse paths tend to have higher goal reward $s_{\text{goal}}$.
@@ 6 Conclusion
First, the current PPR approach is limited to discrete states or sequences. Future work could explore extending PPR to more complex decision-making scenarios that involve continuous states and diverse data modalities. Another limitation is the dependency on the availability of an agent’s policy, which must be either provided or trained. Therefore, future research could investigate the application of PPR with limited training data. However, since there is extensive research on policy estimation from limited data, this aspect is beyond the scope of this paper.
@@ A.1 Grid-world
In this experiment, we create a 8x6 array and set $b=200,k=1,c_{e}=1,\gamma=0.99,\lambda_{\text{policy}}=0.1,\lambda_{\text{path}}=0.1,\text{decay}\>\epsilon=0.001,C=1$ to run the Algorithm 1. The learning rate was $31e-$ 3 We select 10 random original routes and generate 10 corresponding recourse paths. We report one example in Figure 1 in the main paper.
@@
We conduct sensitivity analysis using the same 10 original routes but with different values of $\lambda_{\text{policy}},\lambda_{\text{path}}$.We report the heatmaps of the average scores of $s_{\text{policy}},s_{\text{path}},s_{\text{goal}}$ in Figure 6. In general, the higher values of $\lambda_{\text{policy}}$ and $\lambda_{\text{path}}$ are, the higher $s_{\text{policy}}$ and $s_{\text{path}}$ we can get. When both of these values are low, such as 0.01, the PPRs tend to have higher goal reward $s_{\text{goal}}$.
(a) Personalization $s_{\text{policy}}$
(b) Similarity $s_{\text{path}}$
@@
Figure 6: Personalization, similarity and goal score $s_{\text{policy}},s_{\text{path}},s_{\text{goal}}$ for the grid-world experiment. The x-axis indicates different values of $\lambda_{\text{path}}$ and the y-axis shows different values of $\lambda_{\text{policy}}$.
@@ A.2 Mario Game
In this paper, we implement PPR based on the Mario Reinforcement Learning tutorial from https://pytorch.org/tutorials/intermediate/mario“˙rl“˙tutorial.html by Yuansong Feng, Suraj Subramanian, Howard Wang, and Steven Guo. The tutorial gives instructions on how to train Mario agent to play itself by using Double Deep Q-Networks) using PyTorch. In this tutorial, Mario’s action space is limited to only two actions: Walk Right or Jump Right. The real state is created by stacking up 3 consecutive game frames, which are also reduced to grayscale. For all the experiments, we train Mario on the first stage/mission only.
@@
To implement PPR, we set $\lambda_{\text{path}}=100$ and $\lambda_{\text{route}}=10$. All the policies are trained with at least 50,000 episodes. Figure 7 shows the probability distribution of taking action ”Jump Right” for all policies: coin seeker, coin dodger, and their PPR policies for the first 460 steps (x-coordinate).
@@ A.3 Temperature
To run Algorithm 1 in the main paper, we set $b=500,k=1,\gamma=0.99,c_{e}=1,\lambda_{\text{policy}}=0.1,\lambda_{\text{path}}=10,\text{decay}\>\epsilon=0.001,C=1$. The learning rate value is $31e-$ 3. We train the generator and the classifier using the standard LSTM architecture.
@@
We also conduct sensitivity analysis with values for $\lambda_{\text{policy}},\lambda_{\text{path}}$. Figure 10 shows the heatmaps of the average score of $s_{\text{policy}},s_{\text{path}},s_{\text{goal}}$ based on different values of $\lambda_{\text{policy}},\lambda_{\text{path}}$. Similar to the grid-world experiment, in general, the larger $\lambda_{\text{policy}}$ and $\lambda_{\text{path}}$, the higher $s_{\text{policy}}$ and $s_{\text{path}}$ are. When both of these values are low, such as 0.01, the PPRs tend to have higher goal reward $s_{\text{goal}}$.
(a) Personalization $s_{\text{policy}}$
(b) Similarity $s_{\text{path}}$
@@
Figure 10: Personalization, similarity and goal score $s_{\text{policy}},s_{\text{path}},s_{\text{goal}}$ for the temperature experiment. The x-axis indicates different values of $\lambda_{\text{path}}$ and y-axis shows different values of $\lambda_{\text{policy}}$.
@@ A.4 Text generation
The language models are trained with GPU NVIDIA GeForce GTX 1660 Ti. accelerator in three days. Our transformer has an embedding dimension size of 200, the dimension of the feedforward network model is 200, 2 transformer encoding layers, and 2 multi-head attention. We also set the dropout rate to 0.2. We set $b=500,k=1,\gamma=0.99,c_{e}=0.01,\text{decay}\>\epsilon=0.001,C=1$. In this experiment, we modify values of $\lambda_{\text{route}},\lambda_{\text{path}}$ before generating the recourse text.
@@
Original text and text sampled from $\mathcal{P}_{A}$ trained on Harry Potter and Bible corpus (with the corresponding $\lambda_{\text{path}}$ value)
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{path}}=10$: I am sorry, it is my fault. $\lambda_{\text{path}}=0.1$: ”I didn’t. Feel so stupid”, said Ron. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{path}}=10:$ I am in distress. $\lambda_{\text{path}}=0.1:$ I have seen my face and my sorrow is stirred.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{path}}=10$: The world is very nosy. $\lambda_{\text{path}}=0.1$: The world and the cup is not strong. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{path}}=10$: The world is fallen. $\lambda_{\text{path}}=0.1$: The world were crucified with him but saved us alive.
@@
$\mathcal{P}_{A}$ trained from Harry Potter corpus: $\lambda_{\text{path}}=10$: The magician had not committed this crime. $\lambda_{\text{path}}=0.1$: The magic was expelled at Hogwarts. $\mathcal{P}_{A}$ trained from Bible corpus: $\lambda_{\text{path}}=10$: The magicians of Egypt as an adversary and did as an evil man. $\lambda_{\text{path}}=0.1$: The magicians of Egypt, but behold, they will bring their evil against their houses.